A listener at a distance from the source hears a sound with intensity of 7.14 x 10-7 W/m2.  

When the listener moved farther from the source to a distance of 132.5 m, he hears an intensity of 2.94 x 10-7 W/m2. What is its initial distance from the source of sound?


1
Expert's answer
2021-02-17T11:00:10-0500

The intensity of the sound is inversely proportional to the distance from it. Thus:



"I_1 \propto\dfrac{1}{r_1^2}\\
I_2 \propto\dfrac{1}{r_2^2}"

where "I_1 =7.14 \times 10^{-7}W/m^2 , I_2=2.94 \times 10^{-7}W/m^2" are the initial and the final intensity respectively, "r_1,r_2 = 132.5m" are the initial and the final distances. Dividing the second equation by the first one, and expressing "r_1", obtain:


"\dfrac{I_2}{I_1} = \dfrac{r_1^2}{r_2^2} \\
r_1 = r_2\sqrt{\dfrac{I_2}{I_1}}\\
r_1 = 132.5\cdot \sqrt{\dfrac{2.94\times 10^{-7}}{7.14\times 10^{-7}}}\approx 85m"

Answer. 85 m.